OpenStudy (anonymous):

Please HELP! MEDAL! Which of the following is the explicit formula for the sequence below? 3, 11, 19, 27, ... A. an = 3 + 8(n - 1) B. an = 8 + 3n C. an = 3 + 8n D. an = 8 + 3(n - 1)

OpenStudy (anonymous):

first recal the formula for arithmetic sequences: an = a1 + (n-1)*d can u identify what 'd' is? it is the difference from one number to the next.

OpenStudy (anonymous):

:/ :/ :/ Identify D? What do you mean

OpenStudy (anonymous):

'd' is the difference between each number in the sequence of 3, 11, 19, 27

OpenStudy (anonymous):

2, -8, -8

OpenStudy (anonymous):

yep, you get the idea, but the method is like this: d = 11-3 or d = 19-11 so we get positive 8 as 'd' next is to identify a1 of the formula i posted earlier. a1 represents the first number in the sequence, which is..?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

correct so, d = 8 and a1 = 3 substitute those into the formula: an = a1 + (n-1)d